# Learning to Generate Parameters of ConvNets for Unseen Image Data

### v1

@@ Abstract
Typical Convolutional Neural Networks (ConvNets) depend heavily on large amounts of image data and resort to an iterative optimization algorithm (e.g., SGD or Adam) to learn network parameters, which makes training very time- and resource-intensive. In this paper, we propose a new training paradigm and formulate the parameter learning of ConvNets into a prediction task: given a ConvNet architecture, we observe there exists correlations between image datasets and their corresponding optimal network parameters, and explore if we can learn a hyper-mapping between them to capture the relations, such that we can directly predict the parameters of the network for an image dataset never seen during the training phase. To do this, we put forward a new hypernetwork based model, called PudNet, which intends to learn a mapping between datasets and their corresponding network parameters, and then predicts parameters for unseen data with only a single forward propagation. Moreover, our model benefits from a series of adaptive hyper recurrent units sharing weights to capture the dependencies of parameters among different network layers. Extensive experiments demonstrate that our proposed method achieves good efficacy for unseen image datasets on two kinds of settings: Intra-dataset prediction and Inter-dataset prediction. Our PudNet can also well scale up to large-scale datasets, e.g., ImageNet-1K. It takes 8967 GPU seconds to train ResNet-18 on the ImageNet-1K using GC from scratch and obtain a top-5 accuracy of 44.65 %. However, our PudNet costs only 3.89 GPU seconds to predict the network parameters of ResNet-18 achieving comparable performance (44.92 %), more than 2,300 times faster than the traditional training paradigm.
@@
Parameter generation, hypernetwork, adaptive hyper recurrent units.
@@ I Introduction
Convolutional Neural Networks (ConvNets) have yielded superior performance in a variety of fields in the past decade, such as computer vision [1], reinforcement learning [2, 3], etc. One of the keys to success for ConvNets stems from huge amounts of image training data. In order to optimize ConvNets, the traditional training paradigm takes advantage of an iterative optimization algorithm (e.g., SGD) to train the model in a mini-batch manner, leading to huge time and resource consumption. For example, when training ResNet-101 [4] on ImageNet [5], it often takes several days or weeks for the model to be well optimized with GPU involved. Thus, how to accelerate the training process of ConvNets is an emergent topic in deep learning.
Nowadays, many methods for accelerating training of deep neural networks have been proposed [6, 7, 8]. The representative works include optimization based techniques by improving the stochastic gradient descent [6, 9, 10], normalization based techniques [7, 11, 12], parallel training techniques [8, 13], et. Although these methods have showed promising potential to speed up the training of the network, they still follow the traditional iterative-based training paradigm.
In this paper, we investigate a new training paradigm for ConvNets. In contrast to previous works accelerating the training of the network, we formulate the parameter training problem into a prediction task: given a ConvNet architecture, we attempt to learn a hyper-mapping between image datasets and their corresponding optimal network parameters, and then leverage the hyper-mapping to directly predict the network parameters for a new image dataset unseen during training. A basic assumption behind the above prediction task is that there exists correlations between image datasets and their corresponding parameters of a given ConvNet.
@@
In order to demonstrate the rationality of this assumption, we perform the following experiment: for an image dataset, we first randomly sample 3000 images to train a 3-layer convolutional neural network until convergence. Then we conduct the average pooling operation to the original inputs as a vector representation of the training data. We repeat the above experiment 1000 times, and thus obtain 1000 groups of representations and the corresponding network parameters. Finally, we utilize Canonical Correlation Analysis (CCA) [17] to evaluate the correlations between training data and the network parameters by the above 1000 groups of data. Fig. 1 shows the results, which illustrates there are indeed correlations between training datasets and their network parameters for a given network architecture.
In light of this, we propose a new hypernetwork based model, called PudNet, to learn a hyper-mapping between image datasets and ConvNet parameters. Specifically, PudNet first summarizes the characters of image datasets by compressing them into different vectors as their sketch. Then, PudNet extends the traditional hypernetwork [18] to predict network parameters of different layers based on these vectors. Considering that model parameters among different layers should be not independent, we design an adaptive hyper recurrent unit (AHRU) sharing weights to capture the relations among them, thereby improving the performance of PudNet. Finally, it is worth noting that for training PudNet, it is prohibitive if preparing thousands of datasets and training the network of a given architecture on these datasets to obtain the corresponding optimal parameters respectively. Instead, we adopt a meta-learning based approach  [19] to train the hypernetwork.
In the experiment, our PudNet showcases the remarkable efficacy for unseen yet related image datasets. For example, it takes around 54, 119, 140 GPU seconds to train ResNet-18 using Adam from scratch and obtain top-1 accuracies of 99.91%, 74.56%, 71.84% on the Fashion-MNIST [14], CIFAR-100 [15], Mini-ImageNet [16], respectively. While our method costs only around 0.5 GPU seconds to predict the parameters of ResNet-18 and still achieves 96.24%, 73.33%, 71.57% top-1 accuracies on the three datasets respectively, at least 100 times faster than the traditional training paradigm. More surprisingly, PudNet exhibits impressive performance even on cross-domain unseen image datasets, demonstrating good generalization ability of PudNet.
@@
  • We find there are correlations between image datasets and their corresponding parameters of a given ConvNet, and propose a general training paradigm for ConvNets by formulating network training into a prediction task.
  • We extend hypernetwork to learn the correlations between image datasets and the corresponding ConvNet parameters, such that we can directly generate parameters for unseen image data with only a single forward propagation.
  • We design an adaptive hyper recurrent unit, which can take data sketch as input and capture the dependencies of parameters among different network layers, thereby improving the performance of PudNet.
  • We perform extensive experiments on image datasets in terms of both Intra-dataset and Inter-dataset prediction tasks, demonstrating the efficacy of our method. We expect that such results can motivate more researchers to explore along with this research direction.
@@ II-A Hypernetwork
Hypernetwork in [18] aims to decrease the number of training parameters, by training a smaller network to generate the parameters of a larger network on a fixed dataset. Hypernetwork has been gradually applied to various tasks [20, 21, 22, 23, 24, 25, 26, 27]. [22] proposes a task-conditioned hypernetwork to overcome catastrophic forgetting in continual learning. Bayesian hypernetwork [20] is proposed to approximate Bayesian inference in neural networks. [28] and [29] utilize many layer-wise hypernetworks shared by different tasks to generate weights for the adapters [30] of different layers. GHN-2 proposed in [31] attempts to build a mapping between the network architectures and network parameters, where the dataset is always fixed. HyperTransformer [32] adopts a transformer-based hypernetwork to blend the information of a small support set to generate parameters for a small CNN in the few-shot setting. However, due to the high time complexity of the attention operation, this approach faces the scalability issue as networks become deeper and the support set becomes larger. Different from these works, we aim to directly predict the parameters of a given ConvNet for unseen image datasets.
@@ II-B Acceleration of Network Training
Many works have been proposed to speed up the training of deep neural networks, including optimization based methods [33, 6, 9], normalization based methods [7, 12, 11], parallel training methods [8, 13, 34], etc. Optimization based methods mainly aim to improve the stochastic gradient descent. For instance, Adaptive Moment Estimation (Adam) [6] achieves this by combining the strengths of AdaGrad [33] and RMSProp [35]. By comprehensively considering both the first-order moment estimation and the second-order moment estimation of the gradients, Adam [6] enables the utilization of more efficient and adaptive learning rates for each parameter. [9] proposes a gradient centralization method that centralizes gradient vectors to improve the Lipschitzness of the loss function. Normalization based methods intend to design good normalization methods to speed up training. The representative work is batch normalization [7] that can make optimization the landscape smooth and lead to fast convergence [36]. Parallel training methods [34, 37] usually stack multiple hardwares to conduct parallel training, which reduces training time by dispersing calculation amounts to distributed devices. However, these methods still follow the traditional iterative based training paradigm. Different from them, we attempt to explore a new training paradigm, and transform the network training problem into a parameter prediction task.
@@ II-C Meta-Learning
Meta-learning (a.k.a. learning to learn) aims to learn a model from a variety of tasks, such that it can quickly adapt to a new learning task using only a small number of training samples [38]. So far, many meta-learning methods have been proposed [39, 40, 38, 41, 42]. The representative algorithms include optimization-based methods and metric-based methods [38]. Optimization-based meta-learning methods usually train the model for easily finetuning by a small number of gradient steps [19, 43]. Metric-based meta-learning methods attempt to learn to match the training sets with the class centroids and predict the label of training sets with the matched classes [40, 44, 45]. Because of its powerful ability, meta-learning has been widely used for various tasks, including few-shot learning [16, 40, 46, 45], zero-shot learning [47], etc. Different from the methods aforementioned, we propose a new training paradigm of neural networks, where we attempt to directly predict network parameters for an unseen dataset with only a single forward propagation without training on the dataset.
Figure 2: Overview of our PudNet. PudNet compresses each image dataset into a sketch with a fixed size, and then utilizes the designed hypernetwork to generate parameters of a target ConvNet by the sketch. Specifically, PudNet exploits the adaptive hyper recurrent units (AHRU) to process the dataset sketch, capturing the dependencies of parameters among different network layers for predicting the parameter representations. Subsequently, the weight generators are employed to produce the dataset-specific parameters of the target ConvNet.
@@ III-A Preliminaries and Problem Formulation
In contrast to traditional iterative-based training paradigm, we attempt to explore a new training paradigm, and formulate the training of ConvNets into a parameter prediction task. To this end, we propose the following objective function:
@@
where $\mathcal{F}(D_{i},\Omega;H_{\theta})$ denotes a forward propagation of our hypernetwork $H_{\theta}$. The input of the forward propagation is the dataset $D_{i}$ and its output is the predicted parameters of ConvNet $\Omega$ by $H_{\theta}$. Note that the architecture of $\Omega$ is always fixed during training and testing, e.g., ResNet-18. This makes sense because we often apply a representative deep model to data of different domains. Thus, it is obviously meaningful if we can predict the network parameters for unseen image data using an identical network architecture. $\mathcal{M}^{\Omega}=\{\mathcal{M}_{i}^{\Omega}\}_{i=1}^{\mathcal{N}}$ denotes the ground-truth parameter set of ConvNet $\Omega$ corresponding to image datasets $\mathcal{D}^{train}$, where $\mathcal{M}_{i}^{\Omega}$ is the ground-truth parameters for the dataset $D_{i}$. $\mathcal{L}$ is a loss function, measuring the difference between the ground-truth parameters $\mathcal{M}_{i}^{\Omega}$ and the predicted parameters.
The core idea in (1) is to learn a hyper-mapping $H_{\theta}$ between image datasets $\mathcal{D}^{train}$ and the ConvNet parameter set $\mathcal{M}^{\Omega}$, on the basis of our finding that there are correlations between image datasets and the ConvNet parameters, as shown in Fig. 1. However, it is prohibitive if preparing thousands of image datasets $D_{i}$ and training $\Omega$ on $D_{i}$ to obtain the corresponding ground-truth parameters $\mathcal{M}_{i}^{\Omega}$ respectively. To alleviate this problem, we adopt a meta-learning based [16] approach to train the hypernetwork $H_{\theta}$, and propose another objective function as:
@@ III-B Overview of Our Framework
Our goal is to learn a hypernetwork $H_{\theta}$, so as to directly predict the ConvNet parameters for an unseen image dataset by $H_{\theta}$. However, there remains two issues that are not solved: First, the sizes of different $D_{i}$ may be different and the dataset sizes may be large, which makes $H_{\theta}$ hard to be trained; Second, there may be correlations among parameters of different layers in a network. However, how to capture such context relations among parameters has not been fully explored so far.
To this end, we propose a novel framework, PudNet, as shown in Fig. 2, PudNet first introduces a dataset compression module to compress each dataset $D_{i}$ into a small size sketch $\mathbf{s}_{i}$ to summarize the major characteristics of $D_{i}$. Then, our context-aware parameter prediction module takes the sketch $\mathbf{s}_{i}$ as input, and outputs the predicted parameters of the target network, e.g., ResNet-18. In addition, multiple adaptive hyper recurrent units sharing weights are constructed to capture the dependencies of parameters among different layers of the network. Finally, PudNet is optimized in a meta-learning based manner.
@@ III-C Dataset Compression
To solve the issue of different sizes of training datasets, we first compress each image dataset into a sketch with a fixed size. In recent years, many data compression methods have been proposed, such as matrix sketching [48, 49], random projection [50], etc. In principle, these methods can be applied to our data compression module. For simplification, we leverage a neural network to extract a feature vector as the representation of each sample, and then conduct the average pooling operation to generate a final vector as the sketch of the dataset. The sketch $\mathbf{s}_{i}$ for the dataset $D_{i}$ can be calculated as:
@@
where $T_{\phi}(\cdot)$ denotes a feature extractor parameterized by $\phi$, and $|D_{i}|$ is the size of the dataset $D_{i}$. The parameter $\phi$ is jointly trained with PudNet in an end-to-end fashion. In our PudNet, the feature extractor contains three basic blocks, where each basic block consists of a $5\times 5$ convolutional layer, a leakyReLU function and a batch normalization layer. For future work, more efforts could be made to explore more effective solutions to summarize the information of a dataset, e.g. using statistic network [51] to compress datasets.

### v3

@@ Abstract
Typical Convolutional Neural Networks (ConvNets) depend heavily on large amounts of image data and resort to an iterative optimization algorithm (e.g., SGD or Adam) to learn network parameters, making training very time- and resource-intensive. In this paper, we propose a new training paradigm and formulate the parameter learning of ConvNets into a prediction task: considering that there exist correlations between image datasets and their corresponding optimal network parameters of a given ConvNet, we explore if we can learn a hyper-mapping between them to capture the relations, such that we can directly predict the parameters of the network for an image dataset never seen during the training phase. To do this, we put forward a new hypernetwork-based model, called PudNet, which intends to learn a mapping between datasets and their corresponding network parameters, then predicts parameters for unseen data with only a single forward propagation. Moreover, our model benefits from a series of adaptive hyper-recurrent units sharing weights to capture the dependencies of parameters among different network layers. Extensive experiments demonstrate that our proposed method achieves good efficacy for unseen image datasets in two kinds of settings: Intra-dataset prediction and Inter-dataset prediction. Our PudNet can also well scale up to large-scale datasets, e.g., ImageNet-1K. It takes 8,967 GPU seconds to train ResNet-18 on the ImageNet-1K using GC from scratch and obtain a top-5 accuracy of 44.65%. However, our PudNet costs only 3.89 GPU seconds to predict the network parameters of ResNet-18 achieving comparable performance (44.92%), more than 2,300 times faster than the traditional training paradigm.
@@
Parameter generation, hypernetwork, adaptive hyper-recurrent units.
@@ I Introduction
Convolutional Neural Networks (ConvNets) have yielded superior performance in a variety of fields in the past decade, such as computer vision [1], reinforcement learning [2, 3], etc. One of the keys to success for ConvNets stems from huge amounts of image training data. To optimize ConvNets, the traditional training paradigm takes advantage of an iterative optimization algorithm (e.g., SGD) to train the model in a mini-batch manner, leading to huge time and resource consumption. For example, when training ResNet-101 [4] on ImageNet [5], it often takes several days or weeks for the model to be well optimized with GPU involved. Thus, how to accelerate the training process of ConvNets is an emergent topic in deep learning.
Nowadays, many methods for accelerating the training of deep neural networks have been proposed [6, 7, 8]. The representative works include optimization-based techniques by improving the stochastic gradient descent [6, 9, 10], normalization based techniques [7, 11, 12], parallel training techniques [8, 13], et. Although these methods have shown promising potential to speed up network training, they still adhere to the traditional iterative-based training paradigm.
In this paper, we investigate a new training paradigm for ConvNets. In contrast to previous works accelerating the training of the network, we formulate the parameter training problem into a prediction task: given a ConvNet architecture, we attempt to learn a hyper-mapping between image datasets and their corresponding optimal network parameters, and then leverage the hyper-mapping to directly predict the network parameters for a new image dataset unseen during training. A basic assumption behind the above prediction task is the presence of correlations between image datasets and their corresponding parameters of a given ConvNet.
@@
To demonstrate the rationality of this assumption, we perform the following experiment: for an image dataset, we first randomly sample 3000 images to train a 3-layer convolutional neural network until convergence. Then we conduct the average pooling operation to the original inputs as a vector representation of the training data. We repeat the above experiment 1000 times and thus obtain 1000 groups of representations and the corresponding network parameters. Finally, we utilize Canonical Correlation Analysis (CCA) [17] to assess the correlations between the training data and the network parameters by the above 1000 data groups. Fig. 1 shows the results, which clearly illustrate the presence of correlations between the training datasets and their network parameters for a given network architecture.
In light of this, we propose PudNet, a new hypernetwork-based model, to learn a hyper-mapping between image datasets and ConvNet parameters. PudNet first summarizes the characteristics of image datasets by compressing them into different vectors as their sketches. Then, it extends the traditional hypernetwork [18] to predict network parameters for different layers based on these vectors. Considering the interdependency of model parameters across layers, we design an adaptive hyper-recurrent unit (AHRU) sharing weights to capture the relations among them. This design enables the predicted parameters of each layer to adapt to both dataset-specific information and the predicted parameter information from previous layers, thereby enhancing the performance of PudNet. Finally, it is worth noting that training PudNet becomes prohibitive if thousands of datasets need to be prepared and the network of a given architecture has to be trained on each of these datasets to obtain their respective optimal parameters. Instead, we adopt a meta-learning based approach [19] to train PudNet.
In the experiment, our PudNet showcases the remarkable efficacy for unseen yet related image datasets. For example, it takes around 54, 119, and 140 GPU seconds to train ResNet-18 using Adam from scratch and obtain top-1 accuracies of 99.91%, 74.56%, and 71.84% on the Fashion-MNIST [14], CIFAR-100 [15], and Mini-ImageNet [16] datasets respectively. In contrast, our method costs only around 0.5 GPU seconds to predict the parameters of ResNet-18 while still achieving top-1 accuracies of 96.24%, 73.33%, and 71.57% on the same three datasets respectively. This represents a substantial improvement as our method is at least 100 times faster than the traditional training paradigm. More surprisingly, PudNet exhibits impressive performance even on cross-domain unseen image datasets, demonstrating the good generalization ability of PudNet.
@@
  • Considering the correlations between image datasets and their corresponding parameters of a given ConvNet, we propose a general training paradigm for ConvNets by formulating network training into a prediction task.
  • We extend hypernetwork to learn the correlations between image datasets and the corresponding ConvNet parameters, enabling the direct generation of parameters for unseen image data with only a single forward propagation. Furthermore, a theoretical analysis of the hyper-mapping is provided to offer interpretability.
  • We design an adaptive hyper-recurrent unit that empowers the predicted parameters of each layer to encompass both the dataset-specific information and the inter-layer parameter dependencies.
  • We perform extensive experiments on image datasets in terms of both Intra-dataset and Inter-dataset prediction tasks, demonstrating the efficacy of our method. We expect that these results will motivate more researchers to explore this research direction.
@@ II-A Hypernetwork
Hypernetwork in [18] aims to decrease the number of training parameters, by training a smaller network to generate the parameters of a larger network on a fixed dataset. Hypernetwork has been gradually applied to various tasks [20, 21, 22, 23, 24, 25, 26, 27]. [22] proposes a task-conditioned hypernetwork to overcome catastrophic forgetting in continual learning. Bayesian hypernetwork [20] is proposed to approximate Bayesian inference in neural networks. [28] and [29] utilize multiple layer-wise hypernetworks shared across different tasks to generate weights for the adapters [30] in different layers. GHN-2 proposed in [31] attempts to build a mapping between the network architectures and network parameters, where the dataset is always fixed. HyperTransformer [32] adopts a transformer-based hypernetwork to blend the information of a small support set to generate parameters for a small CNN in the few-shot setting. However, due to the high time complexity of the attention operation, this approach faces the scalability issue as networks become deeper and the support set becomes larger. Different from these works, we aim to directly predict the parameters of a given ConvNet for unseen image datasets.
@@ II-B Acceleration of Network Training
Many works have been proposed to speed up the training of deep neural networks, including optimization-based methods [33, 6, 9], normalization based methods [7, 12, 11], parallel training methods [8, 13, 34], etc. Optimization-based methods mainly aim to improve the stochastic gradient descent. For instance, Adaptive Moment Estimation (Adam) [6] achieves this by combining the strengths of AdaGrad [33] and RMSProp [35]. By comprehensively considering both the first-order moment estimation and the second-order moment estimation of the gradients, Adam [6] enables the utilization of more efficient and adaptive learning rates for each parameter. [9] proposes a gradient centralization method that centralizes gradient vectors to improve the Lipschitzness of the loss function. Normalization based methods intend to design good normalization methods to speed up training. Batch normalization [7] is a prominent work on this line, known for its ability to smooth the optimization landscape and achieve fast convergence [36]. Parallel training methods [34, 37] commonly leverage multiple hardware resources to conduct parallel training, thereby reducing training time by distributing computational tasks among distributed devices. However, these methods still follow the traditional iterative based training paradigm. Different from them, we attempt to explore a new training paradigm and transform the network training problem into a parameter prediction task.
@@ II-C Meta-Learning
Meta-learning aims to learn a model from a variety of tasks, such that it can quickly adapt to a new learning task using only a small number of training samples [38]. So far, many meta-learning methods have been proposed [39, 40, 38, 41, 42]. The representative algorithms include optimization-based methods and metric-based methods [38]. Optimization-based meta-learning methods usually train the model for easily finetuning by a small number of gradient steps [19, 43]. Metric-based meta-learning methods attempt to learn to match the training sets with the class centroids and predict the label of training sets with the matched classes [40, 44, 45]. Because of its powerful ability, meta-learning has been widely used for various tasks, including few-shot learning [16, 40, 46, 45], zero-shot learning [47], etc. Different from the above methods, we propose a new training paradigm of neural networks, where we attempt to directly predict network parameters for an unseen dataset with only a single forward propagation without training on the dataset.
Figure 2: Overview of our PudNet. PudNet compresses each image dataset into a sketch with a fixed size and utilizes the designed hypernetwork to generate parameters of a target ConvNet using this sketch. Specifically, PudNet exploits the adaptive hyper-recurrent units (AHRU) to process the dataset sketch, capturing the dependencies of parameters among different network layers for predicting the parameter representations. Subsequently, the weight generators are employed to produce the dataset-specific parameters of the target ConvNet.
@@ III-A Preliminaries and Problem Formulation
In contrast to the traditional training paradigm, we attempt to explore a new training paradigm and formulate the training of ConvNets into a parameter prediction task. To this end, we propose the following objective function:
@@
where $\mathcal{F}(D_{i},\Omega;H_{\theta})$ denotes a forward propagation of our hypernetwork $H_{\theta}$. The input of the forward propagation is the dataset $D_{i}$ and its output is the predicted parameters of ConvNet $\Omega$ by $H_{\theta}$. Note that the architecture of $\Omega$ is consistently fixed during both the training and testing stages, e.g., ResNet-18. This choice is logical as we often utilize a representative deep model for data from different domains. Therefore, it becomes significantly meaningful to be able to predict the network parameters for unseen image data using an identical network architecture. $\mathcal{M}^{\Omega}=\{\mathcal{M}_{i}^{\Omega}\}_{i=1}^{\mathcal{N}}$ denotes the ground-truth parameter set of ConvNet $\Omega$ corresponding to image datasets $\mathcal{D}^{train}$, where $\mathcal{M}_{i}^{\Omega}$ is the ground-truth parameters for the dataset $D_{i}$. $\mathcal{L}$ is a loss function, measuring the difference between the ground-truth parameters $\mathcal{M}_{i}^{\Omega}$ and the predicted parameters.
The core idea in (1) is to learn a hyper-mapping $H_{\theta}$ between image datasets $\mathcal{D}^{train}$ and the ConvNet parameter set $\mathcal{M}^{\Omega}$, based on our finding that there are correlations between image datasets and the ConvNet parameters, as shown in Fig. 1. However, it is prohibitive if preparing thousands of image datasets $D_{i}$ and training $\Omega$ on $D_{i}$ to obtain the corresponding ground-truth parameters $\mathcal{M}_{i}^{\Omega}$ respectively. To alleviate this problem, we adopt a meta-learning based [16] approach to train the hypernetwork $H_{\theta}$, and propose another objective function as:
@@ III-B Overview of Our Framework
Our goal is to learn a hypernetwork $H_{\theta}$ to directly predict the ConvNet parameters for an unseen image dataset. There are two issues that remain unresolved. First, the sizes of different $D_{i}$ dataset may vary and the dataset sizes themselves can be large, which poses challenges in training $H_{\theta}$; Second, there may exist correlations among parameters of different layers within a network. However, how to capture such context relations among parameters has not been fully explored.
To this end, we propose a novel framework called PudNet, illustrated in Fig. 2. PudNet begins with a dataset compression module, which compresses each dataset $D_{i}$ into a concise sketch $\mathbf{s}_{i}$, summarizing the key characteristics of $D_{i}$. Next, our context-aware parameter prediction module takes the sketch $\mathbf{s}_{i}$ as input and predicts the parameters of the target network, such as ResNet-18. Furthermore, multiple adaptive hyper-recurrent units with shared weights are constructed to capture parameter dependencies across different layers of the network. Finally, PudNet is optimized in a meta-learning based manner.
@@ III-C Dataset Compression
To address the issue of varying dataset size, we employ a data compression module to compress each image dataset into a fixed-size sketch. In recent years, many data compression methods have been proposed, including matrix sketching [48, 49], random projection [50], etc. These methods can potentially be applied in our data compression module. For simplification, we utilize a neural network to extract a feature vector as the representation of each sample. Subsequently, we perform an average pooling operation to generate a final vector as the sketch of the dataset. The sketch $\mathbf{s}_{i}$ for dataset $D_{i}$ can be calculated as:
@@
where $T_{\phi}(\cdot)$ denotes a feature extractor parameterized by $\phi$, and $|D_{i}|$ is the size of the dataset $D_{i}$. $\phi$ is jointly trained with PudNet in an end-to-end fashion. In our PudNet, the feature extractor contains three basic blocks, each comprising a $5\times 5$ convolutional layer, a leakyReLU function, and a batch normalization layer. For future work, more efforts could be made to explore more effective solutions for summarizing dataset information. One potential direction is to investigate the use of statistic networks [51] to compress datasets.
Discussion. In this paper, we adopt a series of widely used image datasets to evaluate the effectiveness of our method. As the number of classes increases, our method may encounter information loss during the dataset compression step, especially when the class distributions are imbalanced. In such cases, the dataset sketch may be biased towards the head classes, which have a larger number of samples compared to the other classes. To alleviate this, we employ a simple clustering-based method. Specifically, we first perform clustering on the dataset. Then, during the dataset compression process, we average the cluster centroid features of each class to produce the dataset sketch used for parameter prediction. By doing so, we can prevent the dataset sketch from being dominated by head classes and mitigate the problem of information loss from the tail classes.